2003 11: SuSI2 now running under LINUX 2002 12: IPs have been moved to sciops/observing/ip 2000 11: The dewar head has been replaced and the detectors have been backed for cleaning. The zero-points are unaffected 2000 10: M4 changed 1999 11: M4 changed with its spare (with a clean coating), and its alignment improved (no more vignetting). Until Nov. 1999, the lowest part of the CCDs (~500pix) were slightly vignetted by M4. In this data, this region should not be used for accurate photometry. It was also discovered in Jul. 1999 that the coating of M4 was damaged, resulting in a slightly higher diffusion. Both problems were solved here by replacing M4 and realigning it. 1999 11: In the past, the NTT was affected by a problem of scattered light inside the dome, especially during bright Moon. This was greatly improved by painting the inside walls with a high diffusion paint 1999 08: 6 new narrow band filters have been received: He II, H beta, O III and Cont, H alpha and Cont 1998 09: Dewar changed. 1998 08: Parasite light problem fixed (encoder tapped and adapter fitted with velvet). 1998 02: SuSI2 commissioning 1997 12: SuSI2 installed at NTT